# Amortized Bayesian Workflow

### v2

@@ 4 Discussion
We presented an adaptive Bayesian workflow to combine the rapid speed of amortized inference with the undisputed sampling quality of MCMC. Our amortized workflow enables a fundamental shift in the scale and feasibility of Bayesian inference. Applying traditional MCMC (e.g., ChEES-HMC) within a standard Bayesian workflow to every dataset independently would require approximately 10 days of GPU computation across our experimental suite. In contrast, our amortized workflow completes inference in half a day, achieving speedups ranging from over 5×5$\times$5 × to 120×120$\times$120 × depending on the problem. Crucially, high-quality posterior draws are retained through a cascade of diagnostics and selective escalation to PSIS and MCMC. In conclusion, our workflow efficiently uses resources by (i) applying fast amortized inference when the results are accurate; (ii) refining draws with PSIS when possible; and (iii) amortized initializations of slower but accurate MCMC chains when needed.
@@
A key strength of the proposed workflow lies in its modular structure, which allows practitioners to tailor each component to the specific constraints and objectives of their application. For example, the choice of MCMC sampler in Step 3 is fully interchangeable: alternative algorithms such as slice sampling, ensemble samplers (e.g., emcee [22]), or NUTS can be substituted if the model is non-differentiable, multimodal, or requires richer exploration. In settings where GPU resources are limited, launching many parallel MCMC chains on CPUs offers a practical alternative, making the workflow more accessible for a broader range of users.
@@
In cases where low-latency decisions are essential (e.g., real-time experimental pipelines) and likelihood evaluations are expensive, the workflow can operate in a lightweight mode using amortized inference with out-of-distribution rejection alone (i.e., Step 1 in our workflow). Conversely, in high-stakes applications where accuracy is paramount, analysts can route all amortized draws through PSIS and, if needed, proceed to full MCMC to guarantee statistical robustness.
Limitations and future directions.
Training amortized models can require substantial upfront investment in simulation and optimization. This burden may be exacerbated in high-dimensional or weakly identifiable models, where neural estimators may struggle to approximate complex inverse maps. Moreover, while our diagnostic for the amortized posterior draws in Step 1 is effective and highly efficient in practice, it remains an imperfect proxy for the true posterior approximation error and can occasionally result in the acceptance of poor-quality amortized draws (cf. Figure 5). Future work could explore even more expressive discrepancy measures tailored to the task at hand.
@@
More broadly, the workflow supports a compelling vision of training amortized models once and reusing them across tasks or studies. In such settings, our layered diagnostics and selective escalation are crucial for maintaining reliability. This positions the workflow as a practical bridge between amortized inference and traditional Bayesian rigor, enabling scalable yet trustworthy inference.
Acknowledgments and Disclosure of Funding
@@
We thank Alexander Fengler for the help with PyMC implementation of the drift-diffusion model. CL and LA were supported by the Research Council of Finland (grants number 356498 and 358980 to LA). AV acknowledges the Research Council of Finland Flagship program: Finnish Center for Artificial Intelligence, and Academy of Finland project 340721. STR is supported by the National Science Foundation under Grant No. 2448380. MS and PB acknowledge support of Cyber Valley Project CyVy-RF- 2021-16, the DFG under Germany’s Excellence Strategy – EXC-2075 - 390740016 (the Stuttgart Cluster of Excellence SimTech). MS acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under grant agreements No 951847 (ELISE) and No 101070617 (ELSA), and support from the Aalto Science-IT project. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources.
@@ Appendix A Background
In this section, we give a brief review of simulation-based calibration checking, parameter recovery checking, Pareto-smoothed importance sampling, and ChEES-HMC algorithm.
@@
Simulation-based calibration (SBC; [84, 53]) is a principled technique for assessing the calibration of posterior distributions estimated by Bayesian inference procedures, particularly useful in simulation-based amortized inference settings. SBC is based on the idea that if the posterior $p(\theta\,|\,y)$ is correctly specified, then the rank of the true parameter $\theta_{\star}$ among posterior draws should follow a uniform distribution. Formally, SBC defines a test statistic $f:\Theta\times Y\to\mathbb{R}$ (e.g., a component of $\theta$, or the log-likelihood $p(y\,|\,\theta)$). For each simulated dataset $y^{(j)}$ generated from the joint model $p(\theta,y)$, the test statistic is evaluated at the ground-truth parameter $\theta_{\star}^{(j)}$ and compared to the same statistic evaluated over posterior samples $\{\theta_{s}^{(j)}\}\sim q_{\phi}(\theta\,|\,y^{(j)})$. The rank of $f(\theta_{\star}^{(j)},y^{(j)})$ among $\{f(\theta_{s}^{(j)},y^{(j)})\}$ is recorded. Repeating this process for all simulated datasets yields a distribution of rank statistics, which should be uniform under well-calibrated inference. Deviations from uniformity signal systematic bias (e.g., over/under-dispersion) in the posterior approximation. We use the graphical-based approach proposed by Säilynoja et al. [69] to assess the uniformity of the rank statistics in SBC. This method provides visual diagnostics for identifying systematic biases or miscalibrations in the posterior approximation by plotting the empirical cumulative distribution function (ECDF) and confidence bands. Examples of SBC checking results using this approach are provided in Appendix C.
@@
Parameter recovery is a complementary diagnostic to SBC and provides a direct visualization of posterior approximation in recovering true generative parameters [62, 64]. The idea is to simulate a collection of datasets $\{y^{(j)}\}$ along with their corresponding ground-truth parameters $\{\theta_{\star}^{(j)}\}$ from the joint model, and assess whether the posterior distributions $q_{\phi}(\theta\,|\,y^{(j)})$ effectively recover these known values. In our workflow, we compare the posterior median extracted from each posterior to the corresponding ground-truth values, along with the median absolute deviation to indicate uncertainty. These comparisons are visualized using scatter plots, with correlation coefficients quantifying the strength of recovery. While not a direct measure of posterior calibration or correctness, parameter recovery provides important practical insight into whether the learned inverse mapping from $y$ to $\theta$ is effective. Examples of parameter recovery checking results using this approach are provided in Appendix C.
@@
Pareto-smoothed importance sampling (PSIS; [88]) is a robust method for improving the stability and reliability of importance sampling (IS) estimates. Given a target distribution $p(y\,|\,\theta)\,p(\theta)$ and a proposal distribution $q_{\phi}(\theta)$, with samples $\hat{\theta}_{s}\sim q_{\phi}(\theta)$, PSIS stabilizes the raw importance weights $w_{s}=p(y\,|\,\hat{\theta}_{s})\,p(\hat{\theta}_{s})/q_{\phi}(\hat{\theta}_{s}\,|\,y)$ by modeling the tail behavior of the importance weights. Specifically, the distribution of extreme importance weights can be approximated by a generalized Pareto distribution:
@@
where $k$ is the shape parameter, $u$ is the location parameter and $\sigma$ is the scale parameter. The number of finite fractional moments of the importance weight distribution depends on $k$: a generalized Pareto distribution has $1/k$ finite moments when $k>0$. To stabilize the importance sampling estimate, the extreme importance weights are replaced with well-spaced order statistics drawn from the fitted generalized Pareto distribution, leading to a more stable and efficient IS estimator. The estimated shape parameter $\hat{k}$ serves as a diagnostic for the reliability of the importance sampling estimate. For moderate sample size ($S>2000$), $\hat{k}<0.7$ indicates good reliability.

### v3

@@ 4 Discussion
We presented an adaptive Bayesian workflow to combine the rapid speed of amortized inference with the undisputed sampling quality of MCMC. Our amortized workflow enables a fundamental shift in the scale and feasibility of Bayesian inference. Applying traditional MCMC (e.g., ChEES-HMC) within a standard Bayesian workflow to every dataset independently would require approximately 10 days of GPU computation across our experimental suite. In contrast, our amortized workflow completes inference in half a day, achieving speedups ranging from over $5\times$ to $120\times$ depending on the problem. Crucially, high-quality posterior draws are retained through a cascade of diagnostics and selective escalation to PSIS and MCMC. In conclusion, our workflow efficiently uses resources by (i) applying fast amortized inference when the results are accurate; (ii) refining draws with PSIS when possible; and (iii) amortized initializations of slower but accurate MCMC chains when needed.
@@
A key strength of the proposed workflow lies in its modular structure, which allows practitioners to tailor each component to the specific constraints and objectives of their application. In cases where preliminary analysis or low-latency decision-making is essential (e.g., real-time experimental pipelines) or where likelihood evaluations are computationally expensive, the workflow can operate in a lightweight mode using amortized inference with out-of-distribution rejection alone (i.e., Step 1 in our workflow). Conversely, in high-stakes applications where accuracy is paramount, analysts can enforce escalation of all amortized draws through PSIS and, if needed, proceed to full MCMC to guarantee statistical robustness. The choice of MCMC sampler in Step 3 is also fully interchangeable: alternative algorithms such as slice sampling, ensemble samplers (e.g., emcee; Foreman-Mackey et al., 2013), or NUTS can be substituted if the model is non-differentiable, multimodal, or requires richer exploration.
Furthermore, while our paper focuses on the trade-off between wall-clock inference speed and posterior quality, practical deployments may also involve additional factors, such as inference cost (e.g., monetary expense for GPU/CPU hours) and hardware availability. Consequently, the most suitable workflow variant can differ across settings. For example, when GPU resources are limited, launching parallel MCMC chains on CPUs offers a practical alternative, making the workflow more accessible for a broader range of users.
Applicability, limitations and future directions.
Our proposed workflow targets likelihood-based Bayesian models for which prior predictive simulation and likelihood evaluation are possible. It is most beneficial in repeated-inference regimes (many datasets or frequent re-fits), with moderate effective dimensionality, and when a good amortized estimator can be trained once and subsequently reused. Hence, it is not universally suitable and does not yield inference speedup gains for all Bayesian models. Training amortized models requires upfront investment in optimization and simulation. In our experiments, we found that default neural network hyperparameter settings, such as normalizing flow architectures, summary network configurations, and optimizer settings, generally yield good performance.
However, in more challenging cases, such as the GEV problem, adjustments may be necessary, guided by training-phase diagnostics. The simulation burden can be exacerbated in high-dimensional ($\gtrsim 10$ parameters) or weakly identifiable models, where neural estimators may struggle to approximate complex inverse maps. Alternative amortized inference approaches (see, e.g., Mittal et al., 2025) could be explored in future work to complement simulation-based amortized inference in such scenarios. In settings where likelihood evaluations are particularly expensive, iterative refinements of the amortized estimator on individual datasets (Glöckler et al., 2022) may also be a practical alternative to likelihood-based corrections in Steps 2 and 3.
Moreover, while our diagnostic for the amortized posterior draws in Step 1 is effective and highly efficient in practice, it remains an imperfect proxy for the true posterior approximation error and can occasionally result in the acceptance of poor-quality amortized draws (cf. Figure 5). An additional empirical study in Appendix D shows that (1) the Mahalanobis distance and the posterior quality metrics are positively correlated when OOD datasets are present; (2) however, some low-distance datasets still yield poor metrics, highlighting the limitation that the OOD diagnostic cannot fully guarantee accuracy and motivating enforced escalation to PSIS (Step 2) when higher accuracy is a requirement. Future work could explore even more effective discrepancy measures, potentially tailored to the task at hand.
@@
More broadly, the workflow supports a compelling vision of training amortized models once and reusing them across tasks or studies—a strategy well suited to applications ranging from psychology to computational biology, among others. In such settings, our layered diagnostics and selective escalation are crucial for maintaining reliability and efficiency. This positions the workflow as a practical bridge between amortized inference and traditional Bayesian rigor, enabling scalable yet trustworthy inference.
Acknowledgments
@@
We thank Alexander Fengler for the help with PyMC implementation of the drift-diffusion model. CL and LA were supported by the Research Council of Finland (grants number 356498 and 358980 to LA). AV acknowledges the Research Council of Finland Flagship program: Finnish Center for Artificial Intelligence, and Academy of Finland project 340721. STR is supported by the National Science Foundation under Grant No. 2448380. MS and PB acknowledge support of Cyber Valley Project CyVy-RF- 2021-16, the DFG under Germany’s Excellence Strategy – EXC-2075 – 390740016 (the Stuttgart Cluster of Excellence SimTech). PB additionally acknowledges the support of DFG Collaborative Research Cluster 391 “Spatio-Temporal Statistics for the Transition of Energy and Transport” – 520388526. MS acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under grant agreements No 951847 (ELISE) and No 101070617 (ELSA), and support from the Aalto Science-IT project. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources. The authors also acknowledge the research environment provided by ELLIS Institute Finland.
Appendix
This appendix provides additional details and analyses to complement the main text, included in the following sections:
@@ Appendix A Background
This section provides a concise overview of the diagnostics and algorithms used in our workflow, including simulation-based calibration checking, parameter recovery checking, out-of-distribution diagnostic with Mahalanobis distance, Pareto-smoothed importance sampling, the ChEES-HMC algorithm, and the nested $\widehat{R}$ convergence diagnostic. Pseudocodes are also given for reference.
@@
Simulation-based calibration (SBC; Talts et al., 2018; Modrák et al., 2025) is a principled technique for assessing the calibration of posterior distributions estimated by Bayesian inference procedures, particularly useful in simulation-based amortized inference settings. SBC is based on the idea that if the posterior $p(\theta\,|\,y)$ is correctly specified, then the rank of the true parameter $\theta_{\star}$ among posterior draws should follow a uniform distribution. Formally, SBC defines a test statistic $f:\Theta\times Y\to\mathbb{R}$ (e.g., a component of $\theta$, or the log-likelihood $p(y\,|\,\theta)$). For each simulated dataset $y^{(j)}$ generated from the joint model $p(\theta,y)$, the test statistic is evaluated at the ground-truth parameter $\theta_{\star}^{(j)}$ and compared to the same statistic evaluated over posterior samples $\{\theta_{s}^{(j)}\}\sim q_{\phi}(\theta\,|\,y^{(j)})$. The rank of $f(\theta_{\star}^{(j)},y^{(j)})$ among $\{f(\theta_{s}^{(j)},y^{(j)})\}$ is recorded. Repeating this process for all simulated datasets yields a distribution of rank statistics, which should be uniform under well-calibrated inference. Deviations from uniformity signal systematic bias (e.g., over/under-dispersion) in the posterior approximation. We use the graphical approach proposed by Säilynoja et al. (2022) to assess the uniformity of the rank statistics in SBC. This method provides visual diagnostics for identifying systematic biases or miscalibrations in the posterior approximation by plotting the empirical cumulative distribution function (ECDF) and confidence bands (95%). The pseudocode of SBC is provided in Algorithm 1. Examples of SBC checking results using this approach are provided in Appendix C.
Algorithm 1 SBC diagnostic
1:Joint model $p(\theta,y)$; amortized posterior $q_{\phi}(\theta\,|\,y)$; scalar test function $f$; number of simulated datasets $J$ (e.g., 200); posterior draws $S$ (e.g., 1000)
2:for $j=1,\dots,J$ do
3:  Sample $\theta_{\star}^{(j)}\sim p(\theta)$, $y^{(j)}\sim p(y\,|\,\theta_{\star}^{(j)})$
4:  Draw $\theta_{s}^{(j)}\sim q_{\phi}(\theta\,|\,y^{(j)})$ for $s=1,\dots,S$
5:  Compute $T_{\star}^{(j)}=f(\theta_{\star}^{(j)},y^{(j)})$, $T_{s}^{(j)}=f(\theta_{s}^{(j)},y^{(j)})$
6:  Compute rank $r^{(j)}=\sum_{s=1}^{S}\mathbb{I}[{T_{s}^{(j)}<T_{\star}^{(j)}}]+\mathrm{uniform}(0,\sum_{s=1}^{S}\mathbb{I}[{T_{s}^{(j)}=T_{\star}^{(j)}}])$
7:end for
8:Compare empirical ranks $r^{(j)}$ against the $\mathrm{uniform}(0,S)$ distribution using the graphical approach of Säilynoja et al. (2022) to identify miscalibration patterns. Alternatively, the uniformity test based on the scalar statistic in Eq. 7 of Modrák et al. (2025) can also be applied.
@@
Parameter recovery is a complementary diagnostic to SBC and provides a direct visualization of posterior approximation in recovering true generative parameters (Radev et al., 2020; 2023). The idea is to simulate a collection of datasets $\{y^{(j)}\}$ along with their corresponding ground-truth parameters $\{\theta_{\star}^{(j)}\}$ from the joint model, and assess whether the posterior distributions $q_{\phi}(\theta\,|\,y^{(j)})$ effectively recover these known values. In our workflow, we compare the posterior median extracted from each posterior to the corresponding ground-truth values, along with the median absolute deviation to indicate uncertainty. These comparisons are visualized using scatter plots, with correlation coefficients quantifying the strength of recovery. While not a direct measure of posterior calibration or correctness, parameter recovery provides important practical insight into whether the learned inverse mapping from $y$ to $\theta$ is effective. The pseudocode of the parameter recovery diagnostic is provided in Algorithm 2. Examples of parameter recovery checking results using this approach are provided in Appendix C.
Algorithm 2 Parameter recovery diagnostic
1:Joint model $p(\theta,y)$; amortized posterior $q_{\phi}(\theta\,|\,y)$; number of datasets $J$; posterior draws $S$
2:for $j=1,\dots,J$ do
3:  Sample $\theta_{\star}^{(j)}\sim p(\theta)$, $y^{(j)}\sim p(y\,|\,\theta_{\star}^{(j)})$
4:  Draw $\theta_{s}^{(j)}\sim q_{\phi}(\theta\,|\,y^{(j)})$ for $s=1,\dots,S$
5:  for each parameter component $k$ do
6:   Compute posterior median $\hat{\theta}_{k}^{(j)}=\mathrm{median}_{s}[\theta_{s,k}^{(j)}]$
7:   Optionally compute a dispersion measure (e.g., median absolute deviation) for $\theta_{s,k}^{(j)}$
8:  end for
9:end for
10:For each $k$, plot $\hat{\theta}_{k}^{(j)}$ vs. $\theta_{\star,k}^{(j)}$ and report correlation
OOD diagnostic with Mahalanobis distance.
The out-of-distribution (OOD) diagnostic used in Step 1 tests whether an observed dataset $y_{\text{obs}}$ falls outside the support of the prior predictive distribution (i.e., the training distribution for the amortized estimator). We work with low-dimensional summary statistics $s(y)\in\mathbb{R}^{d}$, which are either learned (e.g., via a summary network $h_{\psi}$) or hand-crafted with domain knowledge. In the latter case, the amortized estimator $q_{\phi}$ must be trained using these same hand-crafted statistics. The pseudocode for computing the Mahalanobis distance and checking whether an observed dataset $y_{\text{obs}}$ is OOD is provided in Algorithm 3.
Algorithm 3 OOD diagnostic with Mahalanobis distance (Step 1)
1:Training datasets $\{y^{(m)}\}_{m=1}^{M}$ (e.g., $M=10000$); summary statistic function $s(\cdot)$; rejection level $\alpha$ (e.g., $\alpha=0.05$)
2:Compute $s^{(m)}=s(y^{(m)})$ for all $m$
3:Compute empirical mean $\mu_{s}=\frac{1}{M}\sum_{m=1}^{M}s^{(m)}$ and covariance $\Sigma_{s}=\frac{1}{M}\sum_{m=1}^{M}(s^{(m)}-\mu_{s})(s^{(m)}-\mu_{s})^{\top}$
4:For each $m$, compute Mahalanobis distance $D_{M}(y^{(m)})=\sqrt{(s^{(m)}-\mu_{s})^{\top}\Sigma_{s}^{-1}(s^{(m)}-\mu_{s})}$
5:Let $T_{\alpha}$ be the empirical $(1-\alpha)$-quantile of $\{D_{M}(y^{(m)})\}_{m=1}^{M}$
6:
7:procedure TestOOD($y_{\mathrm{obs}}$)
8:  Compute $s_{\mathrm{obs}}=s(y_{\mathrm{obs}})$ and $D_{M}(y_{\mathrm{obs}})$
9:  return $\mathbb{I}_{D_{M}(y_{\mathrm{obs}})>T_{\alpha}}$ $\triangleright$ 1 = OOD, 0 = in-distribution
10:end procedure
@@
Pareto-smoothed importance sampling (PSIS; Vehtari et al., 2024) is a robust method for improving the stability and reliability of importance sampling (IS) estimates. Given a target distribution $p(y\,|\,\theta)\,p(\theta)$ and a proposal distribution $q_{\phi}(\theta)$, with samples $\hat{\theta}_{s}\sim q_{\phi}(\theta)$, PSIS stabilizes the raw importance weights $w_{s}=p(y\,|\,\hat{\theta}_{s})\,p(\hat{\theta}_{s})/q_{\phi}(\hat{\theta}_{s}\,|\,y)$ by modeling the tail behavior of the importance weights. Specifically, the distribution of extreme importance weights can be approximated by a generalized Pareto distribution (GPD):
@@
where $k$ is the shape parameter, $u$ is the location parameter and $\sigma$ is the scale parameter. The number of finite fractional moments of the importance weight distribution depends on $k$: a generalized Pareto distribution has $1/k$ finite moments when $k>0$. To stabilize the importance sampling estimate, the extreme importance weights are replaced with well-spaced order statistics drawn from the fitted generalized Pareto distribution, leading to a more stable and efficient IS estimator. The estimated shape parameter $\hat{k}$ serves as a diagnostic for the reliability of the importance sampling estimate. The pseudocode of PSIS is provided in Algorithm 4.
Algorithm 4 PSIS weights and Pareto-$\hat{k}$ diagnostic (Step 2)
1:Observed data $y_{\mathrm{obs}}$; log-likelihood $\log p(y\,|\,\theta)$; log-prior $\log p(\theta)$; amortized posterior $q_{\phi}(\theta\,|\,y)$; draws $\theta_{s}\sim q_{\phi}$, $s=1,\dots,S$
2:for $s=1,\dots,S$ do
3:  $\ell_{s}=\log p(y_{\mathrm{obs}}\,|\,\theta_{s})+\log p(\theta_{s})-\log q_{\phi}(\theta_{s}\,|\,y_{\mathrm{obs}})$
4:end for
[…]
7:Sort $w_{s}$} to obtain $w{(1)}\leq\dots\leq w{(S)}$
8:Choose tail size $M=\lfloor\min(0.2S,3\sqrt{S})\rfloor$ and define tail $w_{(S-M+1)},\dots,w_{(S)}$}
9:Fit a GPD to the $M$ largest importance weights $w_{(S-M+1)},\dots,w_{(S)}$} and obtain shape estimate $\hat{k}$
10:Replace the $M$ largest weights with smoothed values from the fitted GPD to obtain stabilized weights $\tilde{w}_{s}$
11:Normalize: $\bar{w}_{s}=\tilde{w}_{s}/\sum_{r=1}^{S}\tilde{w}_{r}$
12:Use $\theta_{s},\bar{w}_{s}$} as weighted PSIS-corrected posterior draws; treat them as reliable if $\hat{k}\leq\min(1-1/\log_{10}(S),0.7)$
Given the PSIS-stabilized weights $\bar{w}_{s}$} from Algorithm 4, one can either compute weighted Monte Carlo estimates directly (Vehtari et al., 2024) or apply the SIR procedure in Algorithm 5 to obtain approximately unweighted posterior draws for downstream use (e.g., visualization or MCMC initialization).
Algorithm 5 Sampling importance resampling (SIR) using PSIS weights
1:PSIS-corrected weighted sample $\{(\theta_{s},\bar{w}_{s})\}_{s=1}^{S}$; desired number of resampled draws $S^{\prime}$ (e.g., $S^{\prime}=S$)
2:Define a categorical distribution on indices $s\in\{1,\dots,S\}$ with probabilities $\bar{w}_{1},\dots,\bar{w}_{S}$
3:for $j=1,\dots,S^{\prime}$ do
4:  Sample index $I_{j}\sim\mathrm{Categorical}(\bar{w}_{1},\dots,\bar{w}_{S})$ $\triangleright$ Typically with replacement; weighted sampling without replacement is useful for generating unique initializations for MCMC chains
5:  Set $\tilde{\theta}_{j}=\theta_{I_{j}}$
6:end for
7:Return unweighted PSIS-corrected posterior draws $\{\tilde{\theta}_{j}\}_{j=1}^{S^{\prime}}$